# Exploring Multilingual Human Value Concepts in Large Language Models: Is Value Alignment Consistent, Transferable and Controllable across Languages?

### v1

@@ 6 Discussions and Suggestions
Our analysis of cross-lingual concept consistency and transferability indicates that multilinguality, dominated by a minority of languages, tends to induce cross-lingual concept inconsistency and unidirectional cross-lingual concept transfer between the dominant language and others. Such patterns could bring about the unidirectional influence of specific knowledge, culture, and even human values of the dominant language onto others, resulting in a low cultural diversity across languages. [^12] A concrete example of such unidirectional cultural impact in the use of LLMs has been found by Zhang et al. (2023): when prompted to write a cover letter in Chinese, ChatGPT frequently generates content containing expressions like “诚挚地 (Sincerely)” and “致意 (Regards)”, which are rare in Chinese but common in English. In contrast, a balanced multilinguality is likely to foster bidirectional cross-lingual transfer, thereby encouraging diversity in culture and human values across languages.
Drawing from our empirical observations and findings, we prudently consider that the following suggestions might contribute to enhancing the safety and utility of multilingual AI. First, we would like to suggest the inclusion of a limited number of dominant languages in pre-training data as source languages for cross-lingual alignment transfer. However, it is essential to simultaneously avoid an excessive prevalence of these languages (exemplified by LLaMA2’s pre-training data, which comprises about 90% English data) to alleviate excessively monotonous transfer patterns, which could potentially further lead to a lack of cultural diversity and increase the risk of multilingual vulnerability. Furthermore, we encourage a more balanced distribution of non-dominant languages to foster mutual cross-lingual transfer patterns, as observed in BLOOMZ models. [^13] These suggestions are based on our findings, which might be biased by factors that we could not observe.
@@ 7 Conclusion
We have presented a systematic exploration of multilingual concepts embedded in LLMs, focusing specifically on human values. Through our extensive analysis spanning 7 human values, 16 languages, and 3 LLM families, we have obtained many interesting findings. Specifically, we empirically verify the presence of multilingual human value concepts in LLMs and observe that the cross-lingual consistency and transferability of these concepts reflect the multilinguality of the models to be extracted. Furthermore, our experiments on cross-lingual control illuminate the multilingual vulnerability of LLMs, as well as the feasibility of cross-lingual control over value alignment of LLMs. With these findings, we prudently present several suggestions for collecting multilingual pre-training data for advanced multilingual AI.
@@
Our work has two limitations as follows: (1) Our primary experimental data rely on translations yielded by translation engines. However, the noise introduced by these translations has minimal impact on our research findings. Firstly, our research focuses on the existence of multilingual human value concepts in LLMs and their multilinguality, which do not depend on exceptional performance in any specific language. Additionally, we examine across multiple tasks, human values, languages, and LLMs to uncover universal patterns, which contributes to the robustness of our results to a certain degree of noise. (2) Constrained by our budgetary resources, we evaluate the effectiveness of model control in a semi-automated manner. This process involves first manually checking a large number of model responses to establish rules and then applying them for further evaluation. In our future work, we plan to explore higher-quality evaluation methods, such as combining manual assessment with AI assistants.
@@
Appendix A Data Details
Our experiments cover 7 concepts of human values: morality, deontology, utilitarianism, fairness, truthfulness, toxicity and harmfulness. Below we describe the definition of each human value along with the public datasets utilized for them.
Morality
The definition of morality revolves around the intuitive acceptance of moral standards and principles that guide individuals in determining the moral status of an act. This set of commonly accepted moral principles is referred to as commonsense morality. For this human value, we utilized the COMMONSENSE MORALITY subset in ETHICS dataset (Hendrycks et al., 2021), which includes first-person characters’ actions with clear moral implications. In detail, for the same scenario, actions with positive or negative moral judgment are provided. The collection of scenarios includes both short and detailed examples, we only utilized the short ones considering our limited computing resources.
@@ Appendix A Data Details
The human value of deontology is defined as the adherence to a set of rules or constraints to determine whether an act is required, permitted, or forbidden. To explore this concept, we employed the DEONTOLOGY subset in ETHICS dataset (Hendrycks et al., 2021), which encompasses two subtasks: Requests and Roles. Specifically, in the Requests subtask, scenarios are created where one character issues a command or request, and another character responds with purported exemptions, which are judged as reasonable or unreasonable. In the Roles subtask, each role is assigned with reasonable and unreasonable responsibilities. We utilized data from both subtasks for our experiments.
@@
Utilitarianism emphasizes maximizing overall well-being, aiming for a world where every individual experiences the highest possible level of well-being. For this concept, we employed the UTILITARIANISM subset in ETHICS dataset (Hendrycks et al., 2021), where pairs of scenarios labeled as either more pleasant or less pleasant are provided.
@@
The human value of fairness involves the equitable treatment of individuals, irrespective of their demographic characteristics. For this human value, we used the StereoSet dataset (Nadeem et al., 2021), which consists of sentences measuring stereotypical bias across gender, race, religion, and profession. These sentences are split into two classes: intrasentence and intersentence. Specifically, each sentence in the intrasentence class has a fill-in-the-blank structure where the blank can be filled with the a stereotype term, anti-stereotype term or unrelated term. We inserted each of these three terms into the blank to form different complete sentences. In the intersentence class, each sentence containing a target term is followed by three associative sentences representing stereotypical, anti-stereotypical, and unrelated associations. We concatenated the preceding and subsequent three types of sentences to form different complete sentences. We only employed pairs of stereotypical and anti-stereotypical sentences to obtain positive and negative samples for this human value.
@@
The human value of truthfulness involves the accurate representation of facts about the real world. In this context, a statement is considered truthful if it aligns with objective reality, without being influenced by personal beliefs or biases. For this concept, we used the TruthfulQA dataset (Lin et al., 2022), which consists of two tasks: generation and multiple-choice. Specifically, in the generation task, questions are accompanied by correct or incorrect responses. In the multiple-choice task, questions are accompanied by a set of candidate answers, some of which are correct and others incorrect. We concatenated the question and its corresponding correct response or answer as a positive example while the same question with its corresponding incorrect response or answer as a negative example.
@@
Toxicity refers to the presence of harmful or offensive language in text, which can include hate speech, harassment, or other forms of harmful communication. For the concept of toxicity, we utilized REALTOXICITYPROMPTS dataset (Gehman et al., 2020) consisting of naturally occurring prompts sampled from English web text and corresponding toxicity scores. We categorized prompts into non-toxic and toxic ones based on the scores, thereby forming positive and negative pairs.
@@
Harmfulness includes various types of detrimental content such as profanity, graphic depictions, threatening behavior, misinformation, discrimination, cybercrime, and dangerous or illegal suggestions. For this concept, we utilized the AdvBench dataset (Zou et al., 2023b) which contains harmful instructions eliciting LLMs to generate objectionable content. These harmful instructions are further combined with harmless instructions to form negative and positive pairs, as described in the work of Zou et al. (2023a).
@@
Language ISO 639-1 Language Family LLaMA2 Ratio(%) BLOOMZ Ratio(%)
English en Indo-European 89.70 30.04
French fr Indo-European 0.16 12.90
Chinese zh Sino-Tibetan 0.13 16.17
Spanish es Indo-European 0.13 10.85
Portuguese pt Indo-European 0.09 4.91
Vietnamese vi Austro-Asiatic 0.08 2.71
Catalan ca Indo-European 0.04 1.10
Indonesian id Austronesian 0.03 1.24
Japanese ja Japonic 0.10 -
Korean ko Koreanic 0.06 -
Finnish fi Uralic 0.03 -
Hungarian hu Uralic 0.03 -
Tamil ta Dravidian - 0.49
Telugu te Dravidian - 0.19
Swahili sw Niger-Congo - 0.01
Chichewa ny Niger-Congo - 0.00007
Table 4: Language distributions of the 16 selected languages (including English), for LLaMA2-chat and BLOOMZ series. Languages ta, te, sw and ny are not included in the pre-training data of LLaMA2-chat series, and languages ja, ko, fi and hu are not included in the pre-training data of BLOOMZ series.
Appendix B Language Distribution
@@
Based on the language distributions in their pre-training data, we categorize the multilinguality pattern of these 3 LLM families into 3 groups: English-dominated LLMs (LLaMA2-chat series in our experiments), Chinese & English-dominated LLMs (i.e., Qwen-chat series), and LLMs with balanced multilinguality (i.e., BLOOMZ series).
Appendix C Complete Results of Multilingual Concept Recognition and Extra Discussions
Complete Results
@@ Appendix C Complete Results of Multilingual Concept Recognition and Extra Discussions
Multilingual Performance Reflects Multilinguality
The performance distributions of different models across all languages reflect their multilinguality. Specifically, while all three model families perform best in English, the LLaMA2-chat series exhibits significant performance disparities between English and non-English languages. The Qwen-chat series, while excelling at English, also outperforms other languages in Chinese. In contrast, the BLOOMZ series demonstrates the smallest performance gap between English and non-English, reflecting a more balanced multilinguality.
Appendix D Computing Pearson Correlation Coefficients Considering Differences in Language Resources
@@
Appendix E Complete Results of Cross-Lingual Concept Consistency and Extra Discussions
Complete Results

### v2

@@ 6 Discussions and Suggestions
Drawing our empirical observations and findings, we prudently consider the following suggestions for the configuration of multilingual pre-training data for LLMs, which might contribute to enhancing multilingual AI safety and utility. First, despite the positive effect of dominant languages as sources for cross-lingual alignment transfer (§5.2), it is essential to avoid an excessive prevalence (exemplified by LLaMA2’s pre-training data, which comprises about 90% English data). Our analysis suggests that such excessive dominance can lead to unfair cross-lingual patterns, manifested as inconsistent multilingual representations (§4.3.1), distorted linguistic relationships (§4.3.2), and monotonous transfer patterns (§4.3.3). These tendencies could potentially further amplify the risk of multilingual vulnerability (§5.2) and undermine cultural diversity (Zhang et al., 2023; Cao et al., 2023). Furthermore, we encourage a more balanced distribution of non-dominant languages, particularly those with extremely limited resources, to foster more equitable cross-lingual patterns (§4.3.1 and §4.3.3). [^6] These suggestions are based on our findings, which might be biased by factors like variations in language performance (§3.4) and other unobserved ones.
@@ 7 Conclusion
We have presented a systematic exploration of multilingual concepts embedded in LLMs, focusing specifically on human value-related concepts (i.e., value concepts). Through our extensive analysis spanning 7 human values, 16 languages, and 3 LLM families, we have obtained many interesting findings. Specifically, we empirically verify the presence of multilingual value concepts in LLMs and identify the cross-lingual characteristics of these concepts arising from language resource disparities. Furthermore, our experiments on cross-lingual control illuminate the multilingual vulnerability of LLMs, as well as the feasibility of cross-lingual manipulation over value alignment of LLMs. With these findings, we prudently present several suggestions for collecting multilingual pre-training data for advanced multilingual AI.
@@
Our work’s major limitation lies in the reliance on translations generated by machine translation for our primary experimental data. A straightforward translation of data related to human values not only introduces translation noise but also overlooks cultural differences. We discuss these two points below.
(1) The noise introduced by machine translations has minimal impact on our research findings. Firstly, our research focuses on the existence of multilingual value concepts in LLMs and their multilinguality, which do not depend on exceptional performance in any specific language. Additionally, we examine across multiple tasks, human values, languages, and LLMs to uncover universal patterns, which contributes to the robustness of our results to a certain degree of noise.
(2) We recognize that cultural variations can result in diverse interpretations of explored values among individuals from different cultural backgrounds. However, our work delves into research questions beyond cultural differences. We primarily focus on the multilingual representations of value concepts with LLMs, their universal cross-lingual patterns, and cross-lingual control over value alignment, aiming to enhance the safety and utility of multilingual AI. Additionally, our proposed framework may also be valuable for studying value disparities. For instance, when applying English concept vectors to other languages for cross-lingual concept recognition, errors in recognition may arise from value disparities between them. We plan to further explore the application of our framework to cultural divergences in our future research.
@@
Appendix A Introduction to the Explored Values
Given that the concepts we delve into are inherently rooted in ethics and morals, it’s essential to clarify their ethical foundations. Below, we present the ethical theory as summarized by Vida et al. (2023). Grounded in this theoretical framework, we then elucidate the definitions and ethical characteristics of each value we explore.
A.1 Ethical Theory
According to Vida et al. (2023), Ethics is divided into four branches: normative ethics, applied ethics, descriptive ethics, and metaethics.
Specifically, normative ethics focuses on the principles and criteria that define moral correctness. It operates within a framework of universal norms and values, providing justification for what is deemed right or wrong. Descriptive ethics, conversely, involves empirical investigations to describe or explain the moral judgments, preferences, and value systems prevalent in societies. It refrains from making moral judgments, focusing instead on documenting and analyzing prevailing ethical beliefs and behaviors. Applied ethics extends the general norms and values from normative ethics to specific contexts and fields, dealing with concrete ethical dilemmas and decisions in domains like bioethics, environmental ethics, or, as relevant to our paper, the ethics of artificial intelligence. Metaethics lays the analytical foundation for these three branches, delving into the nature of moral language, the meaning of moral judgments, and the foundational aspects of ethical theories.
Furthermore, normative ethics can be assigned to three competing ethical families: virtue ethics, deontological ethics, and consequentialism. While deontological ethics emphasizes the intrinsic rightness or wrongness of actions based on principles or rules, consequentialism assesses actions by their outcomes or consequences. Meanwhile, virtue ethics focuses on the moral character and virtues of the individual.
A.2 Definitions and Ethical Characteristics of Each Value
Below, we detail the definitions of the 7 explored values, their ethical characteristics, and any interconnections between them.
Commonsense Morality
Commonsense Morality refers to the intuitive and widely accepted moral principles guiding everyday human behavior. These principles often stem from societal norms, cultural values, and emotional responses, forming the basis of our ethical decision-making. Commonsense Morality focuses on evaluating actions based on moral correctness rather than merely describing existing moral beliefs and behaviors in society. Thus, it can be categorized as a part of normative ethics.
@@ A.2 Definitions and Ethical Characteristics of Each Value
Deontology, on the other hand, focuses on the inherent rightness or wrongness of actions based on adherence to a set of rules or constraints. It asserts that certain actions possess moral obligations or prohibitions, independent of their outcomes. Thus, Deontology is categorized under normative ethics, specifically within the deontological ethics family. While both Commonsense Morality and Deontology belong to normative ethics, they differ in their foundational principles. Commonsense Morality is anchored in societal norms and moral correctness, emphasizing the alignment of actions with shared societal values. In contrast, Deontology prioritizes rule-based morality, focusing on the inherent moral obligations or prohibitions associated with actions, regardless of their outcomes.
@@
Utilitarianism emphasizes maximizing overall well-being, aiming for a world where every individual experiences the highest possible level of well-being. Belonging to the consequentialism family within normative ethics, utilitarianism assesses the moral value of an action based on its outcomes or consequences, contrasting with deontology’s focus on the intrinsic rightness or wrongness of actions.
@@
Fairness pertains to the equitable and impartial treatment of individuals, regardless of their demographic attributes such as race, gender, age, religion, or socioeconomic status. Its emphasis on societal biases places Fairness within the realm of descriptive ethics, focusing less on absolute moral rightness or wrongness.
@@
Truthfulness involves the accurate representation of facts about the real world. In this context, a statement is considered truthful if it aligns with objective reality, without being influenced by personal beliefs or biases. Given that ensuring the honesty and transparency of AI systems is crucial in the realm of artificial intelligence, Truthfulness is more appropriately classified under applied ethics.
@@
Toxicity refers to the presence of harmful or offensive language in text, which can include hate speech, harassment, or other forms of harmful communication. In the context of AI-generated content, Toxicity appropriately falls under applied ethics due to its direct influence on user experience.
@@
Harmfulness includes various types of detrimental content such as profanity, graphic depictions, threatening behavior, misinformation, discrimination, cybercrime, and dangerous or illegal suggestions. Harmfulness is inherently a broader concept and may intersect with other ones. Given its pivotal role in AI alignment research, we classify Harmfulness under applied ethics.
Table 3 further presents the positive and negative examples of each human value. Given the diverse definitions and ethical nature of the concepts we explore, we collectively term them “value concepts” in this paper, also aligning with AI alignment research (Bai et al., 2022; Askell et al., 2021; Hendrycks et al., 2021). Note that the above classification adheres to ethical theories as closely as possible, but some deviation may still exist.
Appendix B Data Details
Below we describe the public datasets utilized for each human value.
Commonsense Morality
We utilized the COMMONSENSE MORALITY subset in ETHICS dataset (Hendrycks et al., 2021), which includes first-person characters’ actions with clear moral implications. In detail, for the same scenario, actions with positive or negative moral judgment are provided. The collection of scenarios includes both short and detailed examples, we only utilized the short ones considering our limited computing resources.
Deontology
We employed the DEONTOLOGY subset in ETHICS dataset (Hendrycks et al., 2021), which encompasses two subtasks: Requests and Roles. Specifically, in the Requests subtask, scenarios are created where one character issues a command or request, and another character responds with purported exemptions, which are judged as reasonable or unreasonable. In the Roles subtask, each role is assigned with reasonable and unreasonable responsibilities. We utilized data from both subtasks for our experiments.
Utilitarianism
We employed the UTILITARIANISM subset in ETHICS dataset (Hendrycks et al., 2021), where pairs of scenarios labeled as either more pleasant or less pleasant are provided.
Fairness
We used the StereoSet dataset (Nadeem et al., 2021), which consists of sentences measuring stereotypical bias across gender, race, religion, and profession. These sentences are split into two classes: intrasentence and intersentence. Specifically, each sentence in the intrasentence class has a fill-in-the-blank structure where the blank can be filled with the a stereotype term, anti-stereotype term or unrelated term. We inserted each of these three terms into the blank to form different complete sentences. In the intersentence class, each sentence containing a target term is followed by three associative sentences representing stereotypical, anti-stereotypical, and unrelated associations. We concatenated the preceding and subsequent three types of sentences to form different complete sentences. We only employed pairs of stereotypical and anti-stereotypical sentences to obtain positive and negative samples for this human value.
Truthfulness
We used the TruthfulQA dataset (Lin et al., 2022), which consists of two tasks: generation and multiple-choice. Specifically, in the generation task, questions are accompanied by correct or incorrect responses. In the multiple-choice task, questions are accompanied by a set of candidate answers, some of which are correct and others incorrect. We concatenated the question and its corresponding correct response or answer as a positive example while the same question with its corresponding incorrect response or answer as a negative example.
Toxicity
We utilized REALTOXICITYPROMPTS dataset (Gehman et al., 2020) consisting of naturally occurring prompts sampled from English web text and corresponding toxicity scores. We categorized prompts into non-toxic and toxic ones based on the scores, thereby forming positive and negative pairs.
Harmfulness
We utilized the AdvBench dataset (Zou et al., 2023b) which contains harmful instructions eliciting LLMs to generate objectionable content. These harmful instructions are further combined with harmless instructions to form negative and positive pairs, as described in the work of Zou et al. (2023a).
@@ Appendix B Data Details
Appendix C Impact of Translation Quality
Our primary experimental data rely on translations yielded by translation engines. However, the noise introduced by these translations has minimal impact on our research findings. Our exploration of universal cross-lingual characteristics in LLMs, such as cross-lingual consistency and transferability, suggests that overall patterns are likely preserved when similar noise affects all languages simultaneously. For example, despite the “translationese effect” which could potentially enhance the similarity between non-English texts and English, significant cross-lingual inconsistencies remain between English and other languages in the LLaMA2-chat-7B series, as illustrated in Figure 2.
en fr zh es pt vi ca id ja ko fi hu Avg
mean 97.5 90.2 91.0 91.7 92.0 84.9 90.2 86.4 87.4 82.7 83.4 81.4 88.2
pca 96.7 92.7 90.7 91.7 89.2 85.9 90.2 83.2 86.9 80.7 82.2 81.2 87.6
Table 5: Comparison of multilingual concept recognition accuracy between PCA-based and mean-based concept extraction methods.
Figure 4: Multilingual concept recognition accuracy across different model layers. Results are averaged across languages included both in LLaMA2-chat and BLOOMZ series’ pre-training data, as well as across all human values.
Figure 5: English concept recognition accuracy with varying numbers of training samples for collecting concept vectors. The result are based on LLaMA2-chat-13B. We calculate the average accuracy across all layers to ensure the results of different settings are comparable.
Appendix D Language Distribution
@@
Based on the language distributions in their pre-training data, we categorize the multilinguality of these 3 LLM families into 3 groups: English-dominated LLMs (LLaMA2-chat series in our experiments), Chinese & English-dominated LLMs (i.e., Qwen-chat series), and LLMs with balanced multilinguality (i.e., BLOOMZ series).
Appendix E More Results of Multilingual Concept Recognition
E.1 Extracting Concept Vectors based on PCA
To further enhance the robustness of our results, we also employed the PCA-based method and compared it with the mean-based approach outlined in Section 3.1 (refer to Hämmerl et al. (2023) or Zou et al. (2023a) for details on the PCA-based method). Table 5 presents the multilingual concept recognition accuracy (Section 3.2) for the concept of deontology on LLaMA2-chat-7B. The results suggest that the mean-based method extracts more distinct concept vectors across languages compared to the PCA-based method, consistent with the conclusions of Zou et al. (2023a).
E.2 Results across Model Layers
Figure 4 presents the concept recognition accuracy across different model layers. We observe that the most explicit concept vectors are encoded in the middle layers of the models. This suggests that the middle layers are more likely to encode explicit abstract value information, aligning with the findings of Li et al. (2023); Zou et al. (2023a).
E.3 Varying the Size of $\mathcal{T}_{c}^{\textnormal{train}}$
We employed varying amounts of training samples to extract concept vectors, and the recognition performance for each human value is illustrated in Figure 5. Surprisingly, optimal accuracy can be achieved for all human values even with few training samples, consistent with the findings by Li et al. (2023), suggesting that the concept vectors for human values are readily extractable in LLMs. Furthermore, we observe notable differences in the recognition accuracy of different human values, indicating different degrees of difficulty in capturing them. Specifically, harmfulness, toxicity, commonsense morality, and deontology are relatively explicitly encoded human values. In contrast, LLMs encounter a greater challenge in recognizing concepts like truthfulness, fairness and utilitarianism.
E.4 Complete Results
@@
E.5 Multilingual Performance Reflects Multilinguality
As shown in Figure 1, the performance distributions of different models across all languages reflect their multilinguality. Specifically, while all three model families perform best in English, the LLaMA2-chat series exhibits significant performance disparities between English and non-English languages. The Qwen-chat series, while excelling at English, also outperforms other languages in Chinese. In contrast, the BLOOMZ series demonstrates the smallest performance gap between English and non-English, reflecting a more balanced multilinguality.
Appendix F Computing Pearson Correlation Coefficients Considering Differences in Language Resources
@@
Appendix G More Results of Cross-Lingual Concept Consistency
same different
LLaMA2-chat-7B (en-en) 1.00 0.56
Qwen-chat-7B (en-en) 1.00 0.49
BLOOMZ-7B1 (en-en) 1.00 0.49
LLaMA2-chat-7B (en-fr) 0.95 0.54
Qwen-chat-7B (en-fr) 0.92 0.44
BLOOMZ-7B1 (en-fr) 0.95 0.53
Table 6: Cosine similarity between concept vectors representing either the same or different values across languages.
G.1 Cosine Similarity between Concept Vectors can Reflect Their Correlation
Steck et al. (2024) discussed the limitations and potential issues with using cosine similarity as a measure of semantic similarity, particularly in the context of embeddings learned from linear models. They highlight that cosine similarity can sometimes produce arbitrary and non-unique results, implying that a high average cosine similarity might raise concerns when dealing with unrelated representations.
In our paper, cosine similarity is calculated on concept vectors across different languages to measure their consistency. It is worth recalling that these concept vectors are computed by averaging a set of difference vectors. This averaging process inherently filters out irrelevant information to some extent, thereby mitigating the unpredictable impact on cosine similarity results.
Furthermore, we attempt to evaluate the effectiveness of cosine similarity outcomes in our specific context. Specifically, we compute the cosine similarity between concept vectors of different values in English (e.g., $cosine(\bm{v}_{c1}^{en},\bm{v}_{c2}^{en})$) and cross-lingually between English (en) and French (fr) for both the same (e.g., $cosine(\bm{v}_{c1}^{en},\bm{v}_{c1}^{fr})$) and different (e.g., $cosine(\bm{v}_{c1}^{en},\bm{v}_{c2}^{fr})$) human values. The averaged results presented in Table 6 indicate that, compared to the same human values, the concept representations of unrelated human values exhibit significantly lower cosine similarity. This observation holds true both within a single language and across languages. These findings suggest that, at least in our context, high cosine similarity tends to indicate high relevance, while low cosine similarity often signifies irrelevance to a considerable extent.
G.2 Results across Model Layers
G.3 Complete Results